OpenStudy (anonymous):

A driver traveled at 40 mph for two hours and 30 mph for one-half hour. What was his average rate of speed?

OpenStudy (anonymous):

it would be total distance over total time... 40 mph for two hours means 80 miles traveled in 2 hours 30 mph for 1/2 hour means 15 miles traveled in 1/2 hour total distance is 95 miles total time is 5/2 hours \[speed = \frac{distance}{time} = \frac{95 miles}{\frac{5}{2} hours} = \frac{95 \times 2}{5} = 19 \times 2 = 38 mph\]

OpenStudy (anonymous):

muplity?

OpenStudy (anonymous):

40 X 2 = 80 miles?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

30 miles X 1/2 hour?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

\[distance = speed \times time\]

OpenStudy (anonymous):

do you have any further questions? i got to go

OpenStudy (anonymous):

How did you get 5/2 hours? Please explain to me.

OpenStudy (kropot72):

If you drive for two hours at 40 miles per hour how many miles would you have travelled?

OpenStudy (kropot72):

@Pluto113 Are you there?

OpenStudy (anonymous):

Yes, I am here.

OpenStudy (anonymous):

40 miles x 2 hours = 80 miles, right?

OpenStudy (anonymous):

30 miles x 1/2 hours = 15 miles, right? I am not sure what the answer, 5/2 hours is...

OpenStudy (anonymous):

That means 2 hours + 1/2 hours = 2.5 or 5/2 hours, right? I just want to comperhend myself.

OpenStudy (kropot72):

And if you travel for 1/2 hour at 30 miles per hour you would have travelled 15 miles. 80 miles are travelled in 2 hours and 15 miles are travelled in 1/2 hour. Total distance = 80 + 15 = 95 miles Total time taken = 2 + 1/2 = 2 1/2 hours = 5/2 hours The mixed number 2 1/2 has been converted to a fraction 5/2 Do you understand?